Convert the following rectangular coordinates into polar coordinates (r , theta) so that r<0 and 0<theta <2pi: ( 4, -4 square root of 3)


1
Expert's answer
2022-06-06T12:42:17-0400
"r^2=(4)^2+(-4\sqrt{3})^2=64"

"r<0=>r=-\sqrt{64}=-8"

Quadrant IV


"\tan\theta=\dfrac{-4\sqrt{3}}{4}=-\sqrt{3}"

"\theta=\pi+\tan^{-1}(-\sqrt{3})=\dfrac{2\pi}{3}"

"(r, \theta)=(-8, \dfrac{2\pi}{3})"